A boy moves from point A in distance 20m in the direction North 60° East to arrive at point B in a direction North 30° East for point C,10m away from point B.What is it vector location from point A?

1
Expert's answer
2022-03-01T15:30:38-0500

"AC=\sqrt{AB^2-BC^2}"

"AC=\sqrt{20^2-10^2}=\sqrt{400-100}=10\sqrt3"